OpenStudy (anonymous):

Solve the Proportion. 1/3 = x/12. For those who prefer wording: (1 over 3 = x over 12) No strait answer, looking for steps.

OpenStudy (anonymous):

1/3*12 = x 12/3 = x 4 = x

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

No worries at all.

OpenStudy (studygurl14):

\[\frac{ 1 }{ 3 } = \frac{ x }{ 12 }\] \[1(12) = 3(x) \rightarrow 12 = 3x \rightarrow x = 4\]